OpenStudy (anonymous):

Solve for x: (3/x)-(2/3)=5/6

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

multiply every term by 6x to eliminate the denominators \[\frac{3\times6x}{x} - \frac{ 2\times6x}{3} = \frac{5 \times6x}{6} \] then 18 - 4x = 5x

OpenStudy (anonymous):

3 - 2x/3 = 5x/6 18/6 - 4x/6 = 5x/6 18/6 = (4x + 5x)/6 18/6 = 9x/6 3 = 3/2x 2 = x

hero (hero):

You don't really have to multiply every term by x to solve this

OpenStudy (anonymous):

another question i am having difficulty with is solve for x: 5/x^2 - 2/x = 2/x+1

OpenStudy (anonymous):

the options are x= 1/4 and x = 2 x= -1/4 and x = -2 x= 1/4 and x=-2 x=-1/4 and x=2

hero (hero):

You should have recognized the mistake I made when I used x - 1 instead of x + 1

hero (hero):

\(\frac{5}{x^2} - \frac{2}{x} = \frac{2}{x+1}\) \(\frac{5}{x^2} - \frac{2x}{x^2} = \frac{2}{x+1}\) \(\frac{5 - 2x}{x^2} = \frac{2}{x+1}\) \(2x^2 = (x+1)(5-2x)\) \(2x^2 = -2x^2 + 3x + 5\) \(4x^2 - 3x - 5=0\) Okay, I'm pretty sure that is correct, so now you have to double check to make sure you put the right problem and the right answer choices.

hero (hero):

You must be talking about the first problem

hero (hero):

\(\frac{3}{x} - \frac{2}{3} = \frac{5}{6}\) \(\frac{3}{x} - \frac{4}{6} = \frac{5}{6}\) \(\frac{3}{x} = \frac{5}{6} + \frac{4}{6}\) \(\frac{3}{x} = \frac{9}{6}\) \(18 = 9x\) \(\frac{18}{9} = x\) 2 = x

hero (hero):

You're not communicating with me so I have no idea which set of answer choices the problem you're working on refers to